What is 4/3c - 3/2 = 8? with directions on how to do it please
Answer by stanbon(75887)

What is 4/3c - 3/2 = 8?
-----
Multiply both sides by 6 to get:
8c - 9 = 48
-----
Add 9 to both sides to get:
8c = 57
------
Divide both sides by 8 to get:
c = 56/8 + 1/8
c = 7 1/8
-----------------
Cheers,
Stan H.
-----------